# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ 4.5 Other early feedback processes
Figure 13: Logarithmic IR dust optical depth in a box of sidelength $2~{}{\rm kpc}$ around the most massive galaxy in the fiducial simulation, projected in the face-on direction. We compute the optical depth with no subgrid model (left panel), a solenoidal subgrid turbulence forcing model (middle panel), and compressive subgrid turbulence forcing model (right panel). In the latter two cases, the lower effective optical depth represents the matter-radiation anti-correlation i.e. radiation escapes preferentially through low density channels in the gas. The strong turbulence in MDGs makes the galaxy optically thin to IR photons when we account for matter-radiation anti-correlation.
Menon et al. (2023) find that although radiation pressure feedback cannot regulate star formation in clouds with surface densities $\Sigma_{\rm gas}\gtrsim 10^{3}~{}{\rm M_{\odot}/pc^{2}}$, the impulse of the FUV radiation force can launch outflows in clouds with surface densities $\Sigma_{\rm gas}\gtrsim 10^{5}~{}{\rm M_{\odot}/pc^{2}}$ which can be accelerated by IR radiation pressure. They propose that this mechanism can explain for the molecular outflows observed by Levy et al. (2021) in super-star clusters (SSCs) of NGC 253. It light of our previous arguments, it remains an open question whether this phenomenon could occur in MDGs.
Besides radiation pressure feedback, we also neglect the effect of stellar winds. Standard stellar mass-loss models, which likely overestimate the effect of stellar winds (Smith, 2014), predict that the momentum injected by stellar winds is at most comparable to that of radiation pressure from UV photons. Furthermore, line-driven winds from O-type stars are expected to weaken with decreasing metallicity. Based on these arguments, we do not think that stellar winds are a dominant feedback mechanism in MDGs, although this possibility should be investigated by future work.
@@ 4.6 Beyond the multi-freefall model
Our star formation recipe is based on the MFF model of Federrath & Klessen (2012). This model assumes that all of the gas in a gravitationally unstable gas cloud at the sonic scale forms stars. However, in reality the process of star formation is not 100% efficient even at the sonic scale. Previous observational (André et al., 2010) and theoretical (Matzner & McKee, 2000) work suggest that only $\epsilon_{\rm core}\sim 30-50\%$ of the gas in a collapsing prestellar core ends up in stars.
@@
The dependence of $\sigma_{s}$, the width of the subgrid density PDF, on the turbulent Mach number $\mathcal{M}_{\rm turb}$ (Eq. 12) comes from fitting a functional form to the small-scale turbulence simulations of Padoan & Nordlund (2011). Those simulations consider the steady state in a gas forced by the idealized Ornstein-Uhlenbeck process.
@@
However, ISM turbulence is not a steady state and the forcing from the bulk velocity field and feedback processes is multi-scale and time-variable. Our turbulence forcing series are only a first pass at characterizing this uncertainty. In addition, the lognormal form of the subgrid density PDF is only an approximation (Federrath et al., 2010; Beattie et al., 2022). Deviations from a lognormal shape are more pronounced at high Mach numbers (Federrath & Klessen, 2012), such as those in MDGs. After a star-formation event, there may be a delay before the lognormal PDF can be reestablished by gravitational collapse which the model does not account for.
@@
On top of these issues, the MFF model is fundamentally limited by its use of the density PDF alone to characterize the unresolved density field. The density PDF is a one-point statistical description that ignores spatial correlation. The same density PDF could be produced by numerous dense subsolar gas clumps or a single mass gas clump, which should physically correspond to different SFEs.
@@
In a recent work, Hennebelle et al. (2024) develop a new star formation recipe they call the turbulent support (TS) model to address some of these shortcomings. The TS model substitutes the lognormal PDF in favor of the more complex Castaing-Hopkins PDF, which has been shown to produce better agreement in the high-Mach regime (Castaing, 1996; Hopkins, 2013).
Compared to the lognormal PDF, the Castaing-Hopkins PDF predicts lower amounts of gas at high densities, a consequence of small-scale turbulent pressure support. In addition, the TS model accounts for the spatial correlation of the unresolved density field, taking the power-law index of the mass function of self-gravitating clumps as a parameter. Finally, the TS model computes the freefall and PDF replenishment timescales and assumes star formation occurs on the minimum of the two.
@@
Brucy et al. (2024) compare the predictions of the TS model to detailed turbulence simulations. Surprisingly, they find that the lognormal PDF reproduces the SFE in the high-Mach regime better than the Castaing-Hopkins PDF. It is unclear whether this discrepancy is physical or related to numerical dissipation in the turbulence simulations. For both PDFs, they find that the MFF tends to overpredict the SFE in the high-Mach regime (their Figure 14) in agreement with Hennebelle et al. (2024). In this regime, the turbulent Jeans length becomes larger than the turbulence injection scale. Because the turbulence injection scale is the maximum size at which turbulent structures can form, turbulence can no longer generate gravitationally unstable structures.
@@
These arguments imply that that MFF may overestimate the SFR in MDGs. However, it’s worth noting that our simulations are at an intermediate resolution $\Delta x_{\rm min}\sim 10~{}{\rm pc}$ relative to the resolution of the detailed turbulence simulations of Brucy et al. (2024) at $\Delta x\sim 2~{}{\rm pc}$ and the size of the simulation domain $\Delta x\sim 1000~{}{\rm pc}$ on which they apply the TS model. It’s possible that although we use a more simple star formation recipe, we still capture some of the relevant behavior simply due to our high resolution. Future work should test the TS model at smaller scales to see whether it converges with the MFF model, and if so, at which scale.
@@
Another aspect of star formation we have neglected is the equation of state (EOS), which we assume to be isothermal. Variations in the EOS can lead to changes in the density statistics (Passot & Vázquez-Semadeni, 1998; Li et al., 2003; Audit & Hennebelle, 2010). The simulations of Glover & Mac Low (2007a, b) indicate that the effective polytropic index may be sub-unity (soft EOS) in the density range $n_{\rm H}\sim 10-10^{4}~{}{\rm cm^{-3}}$ and super-unity (stiff EOS) at much higher densities $n_{\rm H}\gtrsim 10^{9}~{}{\rm cm^{-3}}$ where the gas starts to become optically thick (Masunaga & Inutsuka, 2000). In our simulations, the high-density tail of the subgrid density PDF enters this range (Fig. 17).
@@
Our limited resolution requires that we represent stars as star cluster particles, where each particle represents a bound collection of individual stars. However, it is an open question what fraction of stars form in bound clusters. Observations of the local Universe suggest that the fraction is high, but perhaps not 100% (e.g. Lada & Lada, 2003; Porras et al., 2003; Bressert et al., 2010; Winston et al., 2020). Feedback from non-cluster stars might have an outsized impact on the ISM because such stars likely form in marginally star-forming conditions that are more diffuse.
When we form a star particle, the star particle inherits the velocity of the gas in its parent cell. However, the velocity of the star-forming cloud could deviate from this value on the scale of the turbulent velocity dispersion. Therefore, we might underestimate the diffusion of stars throughout the ISM and the fraction of SN explosions in non-star-forming environments.
@@ 4.7 Feedback modeling caveats
In dense regions when the SN cooling radius is unresolved, we model the SN explosion using momentum feedback. Without momentum feedback, the thermal energy injected by the SN would be unphysically radiated away before the snowplow phase of the SN blastwave.
@@
However, by skipping the earlier phases of the SN explosion, we do not account for the effect of this heating on the star-forming cloud. This small-scale heating could allow SNe to more rapidly shut off star formation within a cloud. If we could resolve these scales, the SN delay time might become important to star formation as predicted in the FFB model (Dekel et al., 2023).
@@
Our star formation recipe assumes that the density field has unresolved structure. Ideally, the same unresolved structure should be accounted for in our SN recipe. Specifically, the SN blastwave should sample the density PDF as it expands. Martizzi et al. (2015) simulate SN explosions in inhomogeneous media with a variety of turbulent Mach numbers. They find that as the Mach number increases, the SN explosion deposits its thermal energy and momentum at progressively larger radii (their Fig. 7). This effect increases the effectiveness of SN feedback. Future modeling should produce power law scalings similar to those in Martizzi et al. (2015) which incorporate the turbulent Mach number as a continuous parameter.
We determine the SN rate at each timestep assuming that SNe are uniformly distributed within the time interval between $\tau_{\rm start}$ and $\tau_{\rm end}$. This approach overestimates the SN rate at the beginning of the star cluster’s life, when the most massive and least abundant stars explode, and underestimates the SN rate at later times, when marginally massive and more abundant stars explode. In future work, we should adjust the SN rate over the course of the star cluster lifetime, following (Kimm et al., 2015).
In our model, SNe always occur at the location of the star cluster particle. However, the stellar population represented by a star cluster particle should have some spatial distribution. Weak two-body interactions can produce “walk-away” stars which slowly diffuse out of the cluster, or strong two-body interactions can produce “run-away” stars which are ejected out of the cluster at rapid velocities. Massive stars have short lifetimes, so walk-away massive stars may still go SN close to the cluster center, but run-away massive stars could go SN far from the cluster center. The fraction of massive stars ejected this way is enhanced relative to less massive stars because massive stars tend to concentrate at the centers of clusters where two-body encounters are more common (Oh & Kroupa, 2016).
If the gas density is more diffuse away from the cluster center, then the ejection of massive stars may decrease the mean gas density in which SNe occur, increasing the effectiveness of SN feedback. This effect was demonstrated in the radiation hydrodynamics simulations of Kimm & Cen (2014). However, the ejection of massive stars from the cluster might also decrease the timescale over which early feedback processes can suppress star formation. Detailed studies of the interaction between cluster dynamics and the ISM are needed to determine the ultimate effect on the SFE.
@@
In the dense galactic environments of Cosmic Dawn, it may be possible for the dense shells of a SN blastwave to themselves be gravitationally unstable and trigger additional star formation. Using simple one-dimensional models, Nagakura et al. (2009) found that a triggered star formation can occur from a typical core-collapse SN in gas with density $n_{\rm H}\gtrsim 10^{3}~{}{\rm cm^{-3}}$. In our simulations, a significant fraction of SNe occur in this regime (Fig 8). Therefore, this effect may increase the SFR relative to what we predict. Secondary star formation might even be enhanced when accounting for small scale turbulence, because the conditions inside a SN blastwave are dominated by compressive forcing modes.
@@ 4.8 Miscellaneous considerations
Our simulations do not include a model for Pop III stars. At $z=9$, this is a good approximation, but our simulations begin at $z\sim 100$ and therefore they include the time period where Pop III stars dominate the cosmic star formation history $z\sim 15-20$ (see Klessen & Glover, 2023, for a review). Pop III stars with masses $140~{}M_{\odot}\lesssim m_{*}\lesssim 260~{}M_{\odot}$ are expected to undergo pair-instability SNe, which have energies 100×100$\times$100 × that of more standard core-collapse SN, leading to more effective SN feedback.
@@ 4.9.1 SFR indicators
4.9.2 Turbulence on observations
The strong turbulence in MDGs may influence our interpretation of observations. In Appendix G, we show that matter-radiation anti-correlation decreases the photon mean free path by a factor proportional to $\mathcal{M}_{\rm turb}^{-2}$. This makes the galaxy less opaque to all photons, not just the IR photons discussed in Section 4.5. We refer to this effect as turbulent deattenuation.
If turbulent deattenuation affects all wavelengths equally, than it would be degenerate with the normalization of the galaxy spectrum and may cause spectral energy distribution (SED) fitting codes to overestimate the stellar mass and SFR. However, the degeneracy could be broken if the turbulent Mach number can be measured from observations.
In the local Universe, this can be done with spatially resolved measurements of the velocity dispersion and temperature. For example, Schneider et al. (2013) used the Herschel Space Observatory’s (Pilbratt et al., 2010) high angular resolution measurements of ${}^{12}{\rm CO}~{}1\to 0$ line widths and dust temperatures to estimate the turbulent Mach number in local molecular clouds. The turbulent Mach number can also be measured using spatially resolved column density maps to estimate the fractal dimension of the density field (Beattie et al., 2019).
The measurement becomes more difficult at Cosmic Dawn, where spatially resolved measurements are not possible, even with high angular resolution, unless the galaxy is strongly lensed. In measurements of the velocity dispersion which are not spatially resolved, contributions from turbulence and from the bulk motion of the gas must be separated, which may not always be possible.

### v2

@@ 4.5.1 Radiation pressure feedback
Figure 14: Logarithmic IR dust optical depth in a box of sidelength $2~{}{\rm kpc}$ around the most massive galaxy in the fiducial simulation, projected in the face-on direction. In the left panel, we compute the contribution of each cell as $\rho\kappa_{\rm IR}$. In the other panels, we compute the contribution of each cell using an effective optical depth accounting for the matter-radiation anti-correlation (App. G) for solenoidal (middle panel) and compressive (right panel) turbulence forcing models. We find that despite the low metallicity, the star-forming regions of the galaxy are optically thick to IR. However, accounting for the matter-radiation anti-correlation significantly reduces the optical depth.
4.5.2 Stellar winds
Line photons from massive stars drive stellar winds which serve as an additional early feedback mechanism (Castor et al., 1975; Weaver et al., 1977; Lancaster et al., 2021). The integrated kinetic energy carried by stellar winds over a massive star lifetime is comparable to that delivered by a SN explosion (Castor et al., 1975). The potential of stellar winds to rival SNe as a feedback mechanism has been shown in simulations (Ceverino & Klypin, 2009; Fierlinger et al., 2016).
The low metallicity in MDGs makes launching line-driven winds more difficult. Even when winds are launched, their impact may be weakened by the leakage of hot gas through low-density channels and energy losses due to turbulent mixing (Rogers & Pittard, 2013; Rosen et al., 2014; Lancaster et al., 2021). These effects are enhanced in the turbulent environment of MDGs. For these reasons, we expect that stellar winds are a subdominant feedback mechanism in MDGs. However, detailed simulations of stellar winds in turbulent, low-metallicity environments are needed to test this hypothesis. Given an appropriate subgrid model, stellar wind feedback could be easily included in our simulations.
@@ 4.6 Beyond the multi-freefall model
Our star formation recipe is based on the MFF model of Federrath & Klessen (2012) as implemented in RAMSES by Kretschmer & Teyssier (2020). In this section, we discuss the shortcomings of this approach and compare our recipe to other works with similar prescriptions.
4.6.1 Other star formation recipes
Federrath & Klessen (2012) discuss three variations of the MFF model with different choices for the critical density threshold $s_{\rm crit}$ from (Krumholz & McKee, 2005) (KM), (Padoan & Nordlund, 2011) (PN), and (Hennebelle & Chabrier, 2011) (HC). The KM and PN versions simultaneously fit all of their detailed simulations to within a factor of two over two magnitudes in SFR.
Our MFF model is based on the KM version with a modification by Kretschmer & Teyssier (2020). Other works (Kimm et al., 2017; Trebitsch et al., 2017; Mitchell et al., 2018; Rosdahl et al., 2018) use a MFF model based on the PN version, except that they only allow stars to form in cells larger than the turbulent Jeans length. This thermo-turbulent model differs from our model because it does not allow star formation in unresolved locally Jeans-unstable clouds if they reside in a cell which is globally Jeans-stable. The thermo-turbulent model has been extended to include magnetic pressure support (Martin-Alvarez et al., 2020) and has been incorporated into multi-physics simulations including on-the-fly radiative transfer and cosmic rays (Martin-Alvarez et al., 2023; Dome et al., 2024).
There are a plethora of star formation recipes outside the MFF model framework. Based on detailed MHD simulations, Padoan et al. (2012) propose an exponential SFE law of the form $\epsilon_{\rm ff}\propto\exp(\alpha_{\rm vir}^{1/2})$, noting that the local SFE is nearly independent of turbulent Mach number. More recent simulations by Kim et al. (2021) also favor the exponential model, although with a different numerical coefficient in the argument of the exponential. This model has successfully been applied to measurements of the SFE in the Milky Way (Evans et al., 2022; Elia et al., 2025). Future work is needed to develop a theoretical foundation for the exponential model, and to describe under what conditions it is favored over MFF-type models.
4.6.2 Prestellar core efficiency
We assume that all of the gas in a gravitationally unstable gas cloud at the sonic scale forms stars. However, the process of star formation is not 100% efficient even at the sonic scale. Previous observational (André et al., 2010) and theoretical (Matzner & McKee, 2000) work suggest that only $\epsilon_{\rm core}\sim 30-50\%$ of the gas in a collapsing prestellar core ends up in stars.
@@ 4.6.3 Subgrid density PDF
We assume that the subgrid density PDF is lognormal (Eq. 10) with width given by Equation 12. However, the true small-scale density field may differ significantly from this description. First, Equation 12 comes from fitting a functional form to the small-scale turbulence simulations of Padoan & Nordlund (2011). Those simulations consider the steady state in a gas forced by the idealized Ornstein-Uhlenbeck process. However, ISM turbulence is not a steady state. Forcing from the bulk velocity field and stellar feedback is multi-scale and time-variable.
@@
Second, the lognormal form of the subgrid density PDF is only an approximation (Federrath et al., 2010; Beattie et al., 2022). Deviations from a lognormal shape are more pronounced at high Mach numbers (Federrath & Klessen, 2012), such as those in MDGs. After a star-formation event, there may be a delay before the lognormal PDF can be reestablished by gravitational collapse which the model does not account for.
@@
Third, the MFF model is fundamentally limited by its use of the density PDF alone to characterize the unresolved density field. The density PDF is a one-point statistical description that ignores spatial correlation. The same density PDF could be produced by numerous dense subsolar gas clumps or a single mass gas clump, which should physically correspond to different SFEs.
@@
Recently, Hennebelle et al. (2024) presented the turbulent support (TS) model to address some of these shortcomings. The TS model substitutes the lognormal PDF in favor of the more complex Castaing-Hopkins PDF, which has been shown to produce better agreement in the high-Mach regime (Castaing, 1996; Hopkins, 2013). Compared to the lognormal PDF, the Castaing-Hopkins PDF predicts less gas at high densities, a consequence of small-scale turbulent pressure support. The TS model also accounts for unresolved spatial correlations parameterized by the power-law index of the mass function of self-gravitating clumps.
@@
Brucy et al. (2024) compare the predictions of the TS model to detailed turbulence simulations. Surprisingly, they find that the lognormal PDF reproduces the SFE in the high-Mach regime better than the Castaing-Hopkins PDF. It is unclear whether this discrepancy is physical or related to numerical dissipation in the turbulence simulations. Regardless, they show that the MFF tends to overpredict the SFE in the high-Mach regime (their Figure 14) in agreement with Hennebelle et al. (2024). At high Mach number, the turbulent Jeans length becomes larger than the turbulence injection scale. This implies that turbulence no longer generates gravitationally unstable structures.
@@
These arguments imply that that MFF may overestimate the SFR in MDGs. However, it’s worth noting that our simulations are at an intermediate resolution $\sim 10~{}{\rm pc}$ relative to the resolution of the detailed turbulence simulations of Brucy et al. (2024) at $\sim 2~{}{\rm pc}$ and the size of the simulation domain $\sim 1000~{}{\rm pc}$ on which they apply the TS model. It’s possible that although we use a more simple star formation recipe, we still capture some of the relevant behavior simply due to our high resolution. Future work should investigate whether the TS model converges with the MFF model at smaller scales.
4.6.4 Subgrid equation of state
Another aspect of star formation we have neglected is the subgrid EOS. The detailed turbulence simulations used to calibrate the MFF model assume an isothermal EOS (Padoan & Nordlund, 2011). This is generally appropriate in dense star-forming regions where cooling is efficient. This also matches the behavior of our simulations in dense regions due to the cooling term in the energy equation.
@@ 4.6.4 Subgrid equation of state
Variations in the EOS lead to changes in subgrid density statistics (Passot & Vázquez-Semadeni, 1998; Li et al., 2003; Audit & Hennebelle, 2010). The simulations of Glover & Mac Low (2007a, b) indicate that the effective polytropic index may be sub-unity (soft EOS) in the density range $\sim 10-10^{4}~{}{\rm cm^{-3}}$ and super-unity (stiff EOS) at much higher densities $\gtrsim 10^{9}~{}{\rm cm^{-3}}$ where the gas starts to become optically thick (Masunaga & Inutsuka, 2000). In our simulations, the high-density tail of the subgrid density PDF enters this range (Fig. 18).
@@ 4.7 Feedback modeling caveats
4.7.1 Supernova clustering
Each star cluster particle in our simulations represents a bound collection of individual stars. The star particle mass is tied to the efficiency of photoionization feedback (Sec. 2.5). However, the star particle mass also affects the efficiency of SN feedback through the clustering of SN explosions. A larger star particle mass produces more clustered SN explosions which enhances the efficiency of SN feedback through the formation of superbubbles with delayed radiative cooling (e.g. Sharma et al., 2014; Gentry et al., 2017). However, a smaller star particle mass samples more diffuse environments where the efficiency of SN feedback is higher. Ideally, one could decouple the photoionization feedback model from the star particle mass and sample star particle masses from a cluster mass function parameterized by the local gas conditions. However, although the cluster mass function has some constraints in local Universe (Fall & Chandar, 2012; Adamo et al., 2015), it could differ substantially in MDGs. This is further complicated by the dependence of cluster sizes on the efficiency of early feedback (Smith et al., 2021).
SN clustering is also impacted by the clustering of the star clusters themselves. In our simulations, star particles inherit the velocity of the gas in their parent cell. However, the velocity of the star cluster could deviate from this value on the scale of the turbulent velocity dispersion. This implies that our simulations underestimate the dispersion of star clusters.
Dynamical interactions can also remove stars from the cluster. Weak two-body interactions produce “walk-away” stars which slowly diffuse out of the cluster. Strong two-body interactions produce “run-away” stars ejected at rapid velocities. Run-away stars can travel a significant distance from the cluster center before going SN. The fraction of massive stars ejected this way is enhanced by their tendency to concentrate at the centers of clusters where two-body encounters are more common (Oh & Kroupa, 2016). Hydrodynamics simulations demonstrate that the ejection of massive stars enhances the efficiency of SNe feedback by producing more SN explosions in diffuse regions (Ceverino & Klypin, 2009; Kimm & Cen, 2014).
Finally, only a fraction of stars form in bound clusters, described by the cluster formation efficiency (CFE) (Bastian, 2008). Milky-Way-like galaxies in the local Universe have typical CFEs $\lesssim 15\%$ (Kruijssen, 2012, Tab. 3). Empirically, the CFE increases in higher gas surface density environments, both within and between galaxies (Adamo et al., 2020a), reaching $37\pm 7\%$ in the Milky Way Galactic center (Ginsburg & Kruijssen, 2018) and $30-100\%$ in nearby mergers (Adamo et al., 2020b). Therefore, the CFE is likely higher in MDGs but still $<100\%$. The deviation from $100\%$ may reduce SN clustering and thus the efficiency of SN feedback (Kruijssen, 2012, Sec. 7.3.4).
SN clustering also depends on the distribution of SN in time. We determine the SN rate at each timestep assuming that SNe are uniformly distributed within the time interval between $\tau_{\rm start}$ and $\tau_{\rm end}$. In reality, SN explosions are more clustered towards the end of that interval due to the negative slope of the IMF. Previous work has accounted for this by adjusting the SN rate over the star particle lifetime (Kimm et al., 2015).
4.7.2 Supernova cooling radius
@@
In dense regions when the SN cooling radius is unresolved, we model the SN explosion using momentum feedback. Without momentum feedback, the thermal energy injected by the SN would be unphysically radiated away before the snowplow phase of the SN blastwave. However, by skipping the earlier phases of the SN explosion, we do not account for the effect of this heating on the star-forming cloud. This small-scale heating could allow SNe to more rapidly shut off star formation within a cloud. If we could resolve these scales, the SN delay time might become important to star formation as predicted in the FFB model (Dekel et al., 2023).
@@
The supernova cooling radius should also be affected by the unresolved density field. Martizzi et al. (2015) simulate SN explosions in inhomogeneous media with a variety of turbulent Mach numbers. They find that as the Mach number increases, the SN explosion deposits its thermal energy and momentum at progressively larger radii (their Fig. 7). This effect increases the effectiveness of SN feedback. Future modeling should produce power law scalings similar to those in Martizzi et al. (2015) which incorporate the turbulent Mach number as a continuous parameter.
4.7.3 Secondary star formation
@@
In the dense galactic environments of Cosmic Dawn, it may be possible for the dense shells of a SN blastwave to themselves be gravitationally unstable and trigger additional star formation. Using simple one-dimensional models, Nagakura et al. (2009) found that a triggered star formation can occur from a typical core-collapse SN in gas with density $\gtrsim 10^{3}~{}{\rm cm^{-3}}$. In our simulations, a significant fraction of SNe occur in this regime (Fig 8). Therefore, this effect may increase the SFR relative to what we predict. Secondary star formation might even be enhanced when accounting for small scale turbulence, because the conditions inside a SN blastwave are dominated by compressive forcing modes.
@@ 4.8 Miscellaneous considerations
4.8.1 Pop III stars
We do not include a model for Pop III stars. At $z=9$, this is a good approximation, but our simulations begin at $z\sim 100$ and therefore they include the time period where Pop III stars dominate the cosmic star formation history $z\sim 15-20$ (see Klessen & Glover, 2023, for a review). Pop III stars with masses $140~{}M_{\odot}\lesssim m_{*}\lesssim 260~{}M_{\odot}$ are expected to undergo pair-instability SNe, which have energies 100×100$\times$100 × that of more standard core-collapse SN, leading to more effective SN feedback.
@@ 4.9.1 SFR indicators
Both SFR indicators hide the intrinsic variability in the SFR on short timescales, reducing the observed variability as pointed out by Furlanetto & Mirocha (2022). Following Pallottini & Ferrara (2023), we quantify this effect using the standard deviation of fluctuations in the log SFR $\sigma_{\rm SFR}$, computed using the procedure in Section 3.6. For the raw and JWST f200w SFRs in our fiducial simulation, the standard deviations are $0.28$ and $0.12$. The variability in UV magnitudes is directly related to the variability in the log SFR by $\sigma_{\rm UV}=2.5\sigma_{\rm SFR}$. For the raw and JWST f200w SFRs in our fiducial simulation, this implies UV variabilities $0.83$ and $0.35$ respectively. The latter more accurately characterizes the variability that would actually be observed. This value is significantly smaller than the variability $\sigma_{\rm UV}\gtrsim 1.5$ required to explain the discrepancies between models and data from burstiness alone (Mason et al., 2023; Shen et al., 2023). Our conclusion is similar to Pallottini & Ferrara (2023), expect that we account for the effect of the SFR indicator. However, our simulations do not capture enhanced burstiness due to coupling with radiation and cosmic rays as demonstrated by Dome et al. (2024).